OpenStudy (anonymous):

The product of 4, 5 , and 6 is equal to twice the sum of 10 and what number?

OpenStudy (anonymous):

4*5*6=2(10+x) 10+x=60 x=50

OpenStudy (anonymous):

Oops, I didn't see part of the equation when I wrote my answer. Malhar got it :)